# Concept – An Evaluation Protocol on Conversation Recommender Systems with System-centric and User-centric Factors

### v2

@@ 5 Conclusion
We regard CRS as a social issue, instead of just a technical problem, considering its major goal of persuading users to accept the recommendations. Hence, we revisit current evaluation protocols and pioneer works to provide a comprehensive survey on understanding the impact of CRS’s characteristics in terms of both pros and cons. We then come up with a new evaluation protocol– Concept, which considers both system- and user-centric factors for addressing personalized user needs in particular, with the hope of enhancing user experience in CRS. Such factors include three key characteristics, which are further divided into six primary abilities. In addition, we suggest that even enhanced by "omnipotent" LLMs, CRS lacks practical usability. We pinpoint several significant limitations of current CRS models, including the incapacity to express itself sincerely, a lack of identity-awareness, low reliability towards contextual nuance, and poor coordination for diverse users. However, our current work does not evaluate the attribute-based CRS Lei et al. (2020), as this type of research often ignores the ability to engage in smooth conversations and instead focuses solely on accurately providing recommendations to users within minimal conversation turns. In this case, evaluating attribute-based CRS models seems unfair. To sum up, through Concept, we aim for providing CRS developers and practitioners with a guidance to draw meaning insights from our inclusive and fine-grained evaluation directions.
@@ B.1 Implementation of CRS Models
Statistics Num
# Conversations 6720
Max Turns 10
Avg. Turns 8.92
Persona Types 12
Table 5: Statistics of the generated conversation dataset.
@@ B.2.2 LLM-based User Simulator
Simulation via Prompts. We prompt the simulator with different personas using the persona descriptions generated by ChatGPT (cf. Table 7.). Concept incorporates the Theory of Mind into our simulator to emulate human social cognition Fischer [2023]. This is achieved by prompting the simulator to first assess its current mental state before generating responses, encouraging reflection on its predefined personality traits and social interactions. Refer to Appendix G for prompts. Additionally, there is a maximum of 10 turns allowed in the conversation, and it will only conclude when the simulated user accepts the recommendation. To prevent the simulated users from directly stating the same attribute preferences as the pre-defined values (i.e., the attribute group) when asked about their preferences by the CRS, we include adjectives before each attribute during the simulation, and we allow the user to describe his/her preference using their own words. This approach is reasonable because in real scenarios, users may not use the exact words as the pre-defined attribute values during the conversation. We have summarized the adjusted values for each attribute in Tables 6, achieved by prompting ChatGPT.
@@ B.2.3 LLM-based Evaluator
Based on the findings of previous studies Ye et al. [2023], Wang et al. [2023c, d], we utilize an instance-wise evaluator to conduct detailed assessments. Expanding on earlier research Ye et al. [2023], Wang et al. [2023d], Liu et al. [2023d], we task the evaluator with using fine-grained scoring rubrics to mitigate scoring bias (see Section E for our reliability analysis). The evaluator assigns scores ranging from 1 to 5 based on ability-specific score rubrics, each accompanied by a corresponding description, using the conversation data. For the generation of fine-grained scoring rubrics, we follow the approach of previous works and employ ChatGPT to produce a set of evaluation criteria Saha et al. [2023a]. This set of criteria serves as a starting point for human refinement Li et al. [2023b]. Prior to assigning a score, we require the evaluator to provide a rationale, drawing inspiration from the effectiveness of CoT prompting Wei et al. [2022], Ye et al. [2023]. For implementation details on the evaluator, refer to Appendix G.
@@
Table 7: Persona description generation. We start by converting each combination of personality and age into a single sentence using a template, and then we use ChatGPT to rephrase it into a natural-sounding sentence. Refer to Appendix G for Prompts.
@@ B.2.4 Computational Metrics
Appendix C More Detailed Related Works
CRS Evaluation. The practical effectiveness of CRS has always been questioned Jannach and Manzoor [2020], Jannach et al. [2021]. Addressing this concern necessitates an inclusive evaluation protocol to assess the strengths, weaknesses, and potential risks of CRS. The success of CRS in practice relies primarily on social characteristics, not just technical ones. However, most evaluation protocols primarily focus on superficial and limited evaluation aspects (in isolation) from system-centric perspectives, such as lexical diversity and perplexity of responses Ghazvininejad et al. [2018], Chen et al. [2019], conversational fluency, relevance, informativeness Wang et al. [2022b, a], recommendation effectiveness and efficiency Wang et al. [2023d], Jin et al. [2019], Wärnestål [2005], and rapport building Hayati et al. [2020]. Since they fail to reveal the inadequacy of CRS from user-centric perspectives, these fragmented evaluation protocols are underdeveloped to constitute a standard benchmark and conduct an inclusive evaluation. Recent efforts have attempted to systematically summarize and evaluate the characteristics of CRS, but they have mainly centered on establishing evaluation protocols through person-to-person conversation analysis and questionnaire interviews, lacking quantitative observations on existing CRS models Jin et al. [2023, 2021], Jannach [2022], Siro et al. [2023] and failing to systematically evaluate system-centric characteristics. Refer to Table 8 for better understanding our differences. In summary, existing protocols are considered too superficial and inadequate to uncover the shortcomings of CRS Jannach et al. [2021], Jannach [2023]. To this end, our Concept brings together system- and user-centric perspectives, enabling a more profound and inclusive evaluation protocol and ultimately contributing to the development of more practical CRS Li et al. [2023a].
LLM as User Simulator. Manual evaluation often requires the user to constantly interact with CRS, which are labor-expensive and generally only feasible in an industry lab Huang et al. [2023], Gao et al. [2021]. To this end, LLM-based user simulators continue to be a rapidly evolving research area due to their labor-intensive nature and effectiveness Kocmi and Federmann [2023], Terragni et al. [2023], Tang et al. [2023], Cheng et al. [2023], Afzali et al. [2023]. A recent study on CRS also features interactive evaluation with LLM-based user simulation Wang et al. [2023d], experimentally demonstrating its effectiveness and reliability as a reliable alternative to humans. Taking inspiration from these studies, Concept also harnesses an LLM-based user simulator to interact with the CRS. Moreover, we equip our simulator with elaborately designed instructions and the Theory of Mind to mimic human social cognition Fischer [2023]. This encourages the simulator to reflect on its own predefined personality traits.
LLM as Evaluator. A LLM-based evaluator could be seen as a combination of a LLM and its prompting strategy Zeng et al. [2023], Cohen et al. [2023], Chan et al. [2023], Wang et al. [2023b], Liu et al. [2023a]. Considering its flexibility, the instance-wise evaluator is widely adopted Ye et al. [2023], Wang et al. [2023c], Liu et al. [2023b]. Importantly, previous findings highlight that offering detailed scoring rubrics or criteria contributes to achieving consistent and aligned evaluations with human assessments Liu et al. [2023d]. These findings are further supported by recent work on CRS Wang et al. [2023d], resulting in a dependable LLM-based evaluator as a viable alternative to human evaluators. Inspired by them, we involve ability-specific scoring rubrics to gain reliable and aligned evaluations. We also encourage the evaluator to generate a rationale before assigning a score, further ensuring reliability Ye et al. [2023].
Appendix D Characteristics Categorization
@@ D.3 Personification
Identity. CRS should provide persuasive and honest explanations to boost the acceptance of the recommendations. In real-world situations, users often have little knowledge about the recommended items. A persuasive explanation has the potential to alter a user’s perception of the recommendation, consequently increasing its acceptance Jannach et al. [2021], Zhou et al. [2022], McSherry [2005]. However, generating explanations must be handled with great care to ensure adequacy and avoid resorting to misleading or deceptive strategies (i.e., sales pitches with deceptive tactics), as this could diminish user trust and loyalty towards the CRS Gkika and Lekakos [2014]. More importantly, deceptive strategies also violate the maxim of sincerity required by the cooperation ability (cf. Section D.2). We emphasize the importance of integrating ethical guidelines into CRS and promoting an ethical CRS that mitigates any potential deceptive behavior.
@@
Appendix E Human Evaluation
Previous findings highlight that offering detailed scoring rubrics or criteria contributes to achieving consistent and aligned evaluations with human assessments Liu et al. [2023d], Wang et al. [2023d], ultimately resulting in a dependable LLM-based evaluator as a viable alternative to human evaluators. We also conduct human evaluation to analyze the consistency. Our human evaluation is used to examine the capability of the CHATGPT-based user simulator being a movie seeker, while checking the correlation between the CHATGPT-based evaluation results and human evaluation results. The results are as follows, which indicate the reliability of our Concept.
@@ Appendix E Human Evaluation
Appendix F Additional Analysis
@@
F.1 Overall Performance
We report more results in Figure 9 in terms of the Human Likeness, overall performance, and user satisfaction. This is achieved by prompting the LLM-based evaluator using the detailed results of all ability-specific scores and fine-grained scoring rubrics. Refer Appendix G for details on prompts.
F.2 Fine-grained Analysis on each Benchmark Dataset

### v3

@@ 5 Conclusion
We regard CRS as a social issue, instead of just a technical problem, considering its major goal of persuading users to accept the recommendations. Hence, we revisit current evaluation protocols and pioneer works to provide a comprehensive survey on understanding the impact of CRS’s characteristics in terms of both pros and cons. We then come up with a new evaluation protocol– Concept, which considers both system- and user-centric factors for addressing personalized user needs in particular, with the hope of enhancing user experience in CRS. Such factors include three key characteristics, which are further divided into six primary abilities. In addition, we suggest that even enhanced by "omnipotent" LLMs, CRS lacks practical usability. We pinpoint several significant limitations of current CRS models, including the incapacity to express itself sincerely, a lack of identity-awareness, low reliability towards contextual nuance, and poor coordination for diverse users. With Concept, we provide an overview for researchers as a reference guidance for evaluating CRS and laying the foundation for CRS enhancement.
@@ Appendix A Limitation Analysis of Our Work
Redial OpendialKG
User Types B C K U B C K U
Anticipation 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Boredom 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Confusion 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Curiosity 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Delight 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Disappointment 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Excitement 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Frustration 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Indifference 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Surprise 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Trust 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
Satisfaction 76 76 76 76 64 64 64 64
Age group=Adults 19 19 19 19 16 16 16 16
Age group=Children 19 19 19 19 16 16 16 16
Age group=Seniors 19 19 19 19 16 16 16 16
Age group=Teens 19 19 19 19 16 16 16 16
In Total 912 912 912 912 768 768 768 768
Table 4: Number of conversational attributes. ’B’ stands for BARCOR, ’C’ for CHATCRS, ’K’ for KBRD, ’U’ for UNICRS.
@@ B.2.2 LLM-based User Simulator
Simulation via Prompts. We prompt the simulator with different personas using the persona descriptions generated by ChatGPT (cf. Table 6.). Concept incorporates the Theory of Mind into our simulator to emulate human social cognition Fischer [2023]. This is achieved by prompting the simulator to first assess its current mental state before generating responses, encouraging reflection on its predefined personality traits and social interactions. Refer to Appendix F for prompts. Additionally, there is a maximum of 10 turns allowed in the conversation, and it will only conclude when the simulated user accepts the recommendation. To prevent the simulated users from directly stating the same attribute preferences as the pre-defined values (i.e., the attribute group) when asked about their preferences by the CRS, we include adjectives before each attribute during the simulation, and we allow the user to describe his/her preference using their own words. This approach is reasonable because in real scenarios, users may not use the exact words as the pre-defined attribute values during the conversation. We have summarized the adjusted values for each attribute in Tables 5, achieved by prompting ChatGPT.
@@ B.2.3 LLM-based Evaluator
Based on the findings of previous studies Ye et al. [2023], Wang et al. [2023c, d], we utilize an instance-wise evaluator to conduct detailed assessments. Expanding on earlier research Ye et al. [2023], Wang et al. [2023d], Liu et al. [2023b], we task the evaluator with using fine-grained scoring rubrics to mitigate scoring bias (see Section D for our reliability analysis). The evaluator assigns scores ranging from 1 to 5 based on ability-specific score rubrics, each accompanied by a corresponding description, using the conversation data. For the generation of fine-grained scoring rubrics, we follow the approach of previous works and employ ChatGPT to produce a set of evaluation criteria Saha et al. [2023a]. This set of criteria serves as a starting point for human refinement Li et al. [2023]. Prior to assigning a score, we require the evaluator to provide a rationale, drawing inspiration from the effectiveness of CoT prompting Wei et al. [2022], Ye et al. [2023]. For implementation details on the evaluator, refer to Appendix F.
@@
Table 6: Persona description generation. We start by converting each combination of personality and age into a single sentence using a template, and then we use ChatGPT to rephrase it into a natural-sounding sentence. Refer to Appendix F for Prompts.
@@ B.2.4 Computational Metrics
Appendix C Characteristics Categorization
@@ C.3 Personification
Identity. CRS should provide persuasive and honest explanations to boost the acceptance of the recommendations. In real-world situations, users often have little knowledge about the recommended items. A persuasive explanation has the potential to alter a user’s perception of the recommendation, consequently increasing its acceptance Jannach et al. [2021], Zhou et al. [2022], McSherry [2005]. However, generating explanations must be handled with great care to ensure adequacy and avoid resorting to misleading or deceptive strategies (i.e., sales pitches with deceptive tactics), as this could diminish user trust and loyalty towards the CRS Gkika and Lekakos [2014]. More importantly, deceptive strategies also violate the maxim of sincerity required by the cooperation ability (cf. Section C.2). We emphasize the importance of integrating ethical guidelines into CRS and promoting an ethical CRS that mitigates any potential deceptive behavior.
@@
Appendix D Human Evaluation
Previous findings highlight that offering detailed scoring rubrics or criteria contributes to achieving consistent and aligned evaluations with human assessments Liu et al. [2023b], Wang et al. [2023d], ultimately resulting in a dependable LLM-based evaluator as a viable alternative to human evaluators. We also conduct human evaluation to analyze the consistency. Our human evaluation is used to examine the capability of the CHATGPT-based user simulator being a movie seeker, while checking the correlation between the CHATGPT-based evaluation results and human evaluation results. The results are as follows, which indicate the reliability of our Concept.
@@ Appendix D Human Evaluation
Appendix E Additional Analysis
@@
E.1 Overall Performance
We report more results in Figure 9 in terms of the Human Likeness, overall performance, and user satisfaction. This is achieved by prompting the LLM-based evaluator using the detailed results of all ability-specific scores and fine-grained scoring rubrics. Refer Appendix F for details on prompts.
E.2 Fine-grained Analysis on each Benchmark Dataset
@@ 4 Relevance
Appendix G Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Section 1 and our Abstract.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Appendix A.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] Instead of introducing new CRS models with potential societal ramifications, we assess the practical applications of existing ones.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] We ensure that this paper does not involve any ethical issues.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] No theoretical result is involved
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] No theoretical result is involved
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] https://github.com/huangzichun/Concept4CRS
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We do not involve any model training. However, we offer model evaluation details in Appendix B.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] We present the average results across various age groups of users.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix B.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [N/A] The dataset utilized in this research was dynamically generated in conjunction with our novel evaluation protocol.
    2. (b)
      Did you mention the license of the assets? [N/A]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] https://github.com/huangzichun/Concept4CRS
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A] The dataset utilized in this research was dynamically generated in conjunction with our novel evaluation protocol.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] The dataset utilized in this research was dynamically generated in conjunction with our novel evaluation protocol. There is no personally identifiable information. We carefully prompt the aligned LLMs to avoid any offensive content.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] The dataset utilized in this research was dynamically generated by LLMs
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] The dataset utilized in this research was dynamically generated by LLMs
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] The dataset utilized in this research was dynamically generated by LLMs
Appendix H Datasheet
To clarify, our contribution lies in the evaluation protocol, not the dataset. The dataset is generated dynamically alongside the execution of the protocol. To prevent any confusion regarding the generated dataset, we also included a datasheet for reference. Note that this datasheet adheres to the guidelines outlined in Gebru et al. [2021].
  1. 1.
    Motivation
    1. (a)
      For what purpose was the dataset created?
      This dataset was created to evaluate and analyze the strengths, weaknesses, and potential risks of off-the-shelf CRS models.
    2. (b)
      Who created the dataset and on behalf of which entity?
      This dataset was generated with the aid of LLMs and CRS models, with the primary code developed by Peixin Qin and Chen Huang.
    3. (c)
      Who funded the creation of the dataset?
      [N/A]
    4. (d)
      Any other Comments?
      [No]
  2. 2.
    Composition
    1. (a)
      What do the instances that comprise the dataset represent?
      A conversation between a user simulator and a CRS model.
    2. (b)
      How many instances are there in total?
      There are 6720 conversations in total.
    3. (c)
      Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
      It contains all possible instances.
    4. (d)
      What data does each instance consist of?
      It comprises a multi-turn conversation between a user and a CRS model, along with the user’s persona, age, preferences, and the specific name of CRS model employed.
    5. (e)
      Is there a label or target associated with each instance?
      Each conversation is associated a targeted movie set as the recommendation ground truth.
    6. (f)
      Is any information missing from individual instances?
      [No]
    7. (g)
      Are relationships between individual instances made explicit?
      [No]
    8. (h)
      Are there recommended data splits?
      [No] All gathered data is intended for CRS model evaluation.
    9. (i)
      Are there any errors, sources of noise, or redundancies in the dataset?
      [N/A]
    10. (j)
      Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
      [Yes] We utilize established movie recommendation datasets as the source for user preference candidates.
    11. (k)
      Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)?
      [No]
    12. (l)
      Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
      [No] We dedicate significant effort to reviewing every conversation.
    13. (m)
      Does the dataset relate to people?
      [Yes] We did not involve real human in our dataset. We employed a user simulator.
    14. (n)
      Does the dataset identify any subpopulations (e.g., by age, gender)?
      [Yes] We simulated various users with different age and personas.
    15. (o)
      Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset?
      [No] All users are simulated.
    16. (p)
      Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)?
      [No]
    17. (q)
      Any other comments?
      [No]
  3. 3.
    Collection Process
    1. (a)
      How was the data associated with each instance acquired?
      Every conversation between a user and the CRS is represented as a single instance in our dataset.
    2. (b)
      What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)?
      We resort to the LLMs. See Section 4.2.
    3. (c)
      If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
      [N/A]
    4. (d)
      Who was involved in the data collection process (e.g., students, crowdworkers, contractors), and how were they compensated (e.g., how much were crowdworkers paid)?
      Only authors are involved in the data collection.
    5. (e)
      Over what timeframe was the data collected?
      All data are collected in 2024-01.
    6. (f)
      Were any ethical review processes conducted (e.g., by an institutional review board)?
      [No] No ethical review processes were conducted with respect to the collection and annotation of this data.
    7. (g)
      Does the dataset relate to people?
      [Yes] We did not involve real human in our dataset. We employed a user simulator.
    8. (h)
      Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
      [No] We collected the data by sorting to the LLMs.
    9. (i)
      Were the individuals in question notified about the data collection?
      [N/A]
    10. (j)
      Did the individuals in question consent to the collection and use of their data?
      [N/A]
    11. (k)
      If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses?
      [N/A]
    12. (l)
      Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted?
      [N/A]
    13. (m)
      Any other comments?
      [No]
  4. 4.
    Preprocessing, Cleaning and Labeling
    1. (a)
      Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?
      [No] No preprocessing/cleaning/labeling is needed.
    2. (b)
      Was the "raw" data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)?
      [N/A] The dataset userd in the paper is the raw dataset itself. We release the raw data on the Github.
    3. (c)
      Is the software used to preprocess/clean/label the instances available?
      [N/A]
    4. (d)
      Any other comments?
      [No]
  5. 5.
    Uses
    1. (a)
      Has the dataset been used for any tasks already?
      [No]
    2. (b)
      Is there a repository that links to any or all papers or systems that use the dataset?
      [No]
    3. (c)
      What (other) tasks could the dataset be used for?
      This dataset can help research ways to improve the persuasiveness and honesty of CRS explanations.
    4. (d)
      Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
      [N/A] No preprocessing/cleaning/labeling
    5. (e)
      Are there tasks for which the dataset should not be used?
      The usage of this dataset should be limited to the scope of CRS.
    6. (f)
      Any other comments?
      [No]
  6. 6.
    Distribution
    1. (a)
      Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
      [Yes] We expect other websites re-distribute our dataset.
    2. (b)
      How will the dataset be distributed (e.g., tarball on website, API, GitHub)?
      The dataset could be accessed on a GitHub webpage.
    3. (c)
      When will the dataset be distributed?
      The dataset is already released to the Github.
    4. (d)
      Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
      [N/A] The dataset is generated dynamically alongside the execution of the protocol. No license is needed.
    5. (e)
      Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
      [No]
    6. (f)
      Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
      [No]
    7. (g)
      Any other comments?
      [No]
  7. 7.
    Maintenance
    1. (a)
      Who is supporting/hosting/maintaining the dataset?
      Chen Huang will be responsible for maintaining the dataset.
    2. (b)
      How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
      E-mail addresses are at the top of the paper.
    3. (c)
      Is there an erratum?
      [No] When errors are discerned, we will announce erratum on our website.
    4. (d)
      Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances’)?
      [Yes] We hope to expand this dataset with more CRS models and recommendation scenarios.
    5. (e)
      If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
      [N/A]
    6. (f)
      Will older versions of the dataset continue to be supported/hosted/maintained?
      [Yes]
    7. (g)
      If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
      [No]
    8. (h)
      Any other comments?
      [No]